OpenStudy (anonymous):

A small class has 10students,4 of whom are girls and 6of whom are boys. The teacher is going to choose two of the students at random. What is the probability that the first student chosen will be a girl and the second will be a boy? Write your answer as a fraction in simplest form.

OpenStudy (anonymous):

4/10 * 6/9 = 2/5 * 2/3 = 4/15